# Nature of Andreev bound states in Josephson junctions of triple-point semimetals

### v1

@@ Abstract
We study superconductor-barrier-superconductor (S-B-S) Josephson junctions constructed out of two-dimensional and three-dimensional triple-point semimetals, which feature a threefold degeneracy at a single nodal point. We assume a weak and homogeneous s-wave pairing in each superconducting region, and a potential difference is applied across a piece of normal-state semimetal to create the barrier region. We apply the BdG formalism to compute the wavefunctions of the subgap Andreev bound states (ABSs), considering the thin-barrier limit, in a piecewise-continuous manner. The appropriate boundary conditions at the S-B and B-S junctions allow us to compute the discrete energy eigenvalues ±$\pm|\varepsilon|$ of the ABSs. We get two distinct solutions for $|\varepsilon|$. This result is different from the S-B-S set-ups in graphene and Weyl semimetals (each featuring twofold-degenerate nodal points), where one obtains only one solution for $|\varepsilon|$. Although the three bands of the triple-point fermions comprise a nondispersive flat band, the multifold nature of the bandstructure is responsible for this difference. We also illustrate the behaviour of the Josephson current flowing across the S-B-S junction, contributed by the two pairs of ABSs.
@@ I Introduction
Since the last decade, there has been intensive studies of bandstructures harbouring symmetry-protected band-crossing points in the Brillouin zone (BZ) [1, 2, 3, 4, 5], giving rise to semimetals. The low-energy $\mathbf{k}\cdot\mathbf{p}$ Hamiltonian near the band-crossings give rise to quasiparticles carrying pseudospin values equal to $\varsigma$, when $(2\,\varsigma+1)$ bands touch at the degeneracy point. We note that whereas in high-energy physics, a relativistic electron is described by a spin-1/2 fermion (fixed by Lorentz invariance), the pseudospin of an emergent fermionic excitation (still having the spin-1/2 value) in nonrelativistic condensed matter physics depends on the irreducible representations of the little group of lattice symmetries at high-symmetry points in the BZ for each of the 230 space groups [3]. The dimensionality of the irreducible representation corresponds to the number of bands crossing at the high-symmetry point. In particular, a threefold-degeneracy can arise in a crystal lattice, when the nodal point is protected by (I) symmorphic rotation, combined with mirror symmetries [6, 7, 8], or (II) nonsymmorphic symmetries [3]. A schematic illustration of the triply-degenerate nodal point is shown in Fig. 1(a).
@@
In this paper, we focus on triple-point fermionic qualiparticles carrying preseudospin value $\varsigma=1$, whose effective low-energy continuum Hamiltonian is of the form $\mathbf{k}\cdot\bm{\mathcal{S}}$, where $\bm{\mathcal{S}}$ represents the vector of the three matrices which form the spin-1 representation of the SO(3) group. This leads to the emergence of two-dimensional (2d) and three-dimensional (3d) semimetals with pseudospin-1 quasiparticles [9, 5, 10, 11, 12, 6, 7, 13, 8, 14, 15, 16, 17, 18, 19, 20, 21], sometimes dubbed as “Maxwell fermions” [9], since their pseudospin quantum number is analogous to the spin-1 quantum number of the photons (which are described by the Maxwell equations). The 3d versions with linear isotropic dispersion are the straightforward generalizations of the Weyl semimetal Hamiltonian $\mathbf{k}\cdot\bm{\sigma}$, with $\bm{\sigma}$ representing the vector operator consisting of the three Pauli matrices, thus representing the spin-1/2 representation of the SO(3) group. One can show that the threefold degeneracies carry Chern numbers ±$\pm 2$ [3, 22, 21, 23], analogous to the pseudospin-3/2 Rarita-Schwinger-Weyl (RSW) semimetals [24, 22, 25, 26, 27, 28, 19, 21, 23, 29, 30] harbouring fourfold-degenerate nodal points, thus reflecting the nontrivial topological character of the corresponding bandstructures. Here, we continue the ongoing efforts to unravel the distinct signatures of the triple-point quasiparticles via various transport properties [14, 15, 19, 20, 21]. In particular, the transport property under consideration will be the Josephson current $I_{J}$ [31, 32, 33] arising in configurations consisting of junctions between the normal (abbreviated by “N”) and the superconducting (abbreviated by “S”) phases of 2d and 3d pseudospin-1 semimetals.
@@
When two superconducting regions are coupled to each other by a weak link between them, $I_{J}$ is the equilibrium dissipationless current flowing across the junction. Since the Andreev surface states of the two superconductors hybridize to form Andreev bound states (ABSs) at the junction, they contribute to $I_{J}$. Here, we extend the study of such Josephson effects in 2d and 3d semimetals [34, 35, 36, 37, 38, 39, 30], where the superconducting regions consist of homogeneous s-wave pairing, and the weak link is created by introducing a tunneling barrier. One can consider two alternative arrangements: (I) S-N-S junctions [34, 35]; (II) S-B-S (where “B” indicates a rectangular potential barrier in the N region) junctions [36, 37, 38, 39, 30, 40]. In experiments, superconductivity can be induced in the appropriate region via proximity-effect by placing a conventional s-wave superconductor on top of an electrode made of the semimetal [41], and the barrier regoin can be created by applying a gate voltage $V_{0}$ across N. The schematics of an S-B-S set-up is illustrated in Fig. 1(b), where we demarcate the left superconducting region as “region I”, the middle barrier region as “region II”, and the right superconducting region as “region III”. We will focus on the short-barrier regime, which implies that the barrier thickness $L$ along the propagation direction is taken to be $L\ll\xi$, where $\xi$ is the superconducting coherence length. In other words, $\xi$ denotes the length scale over which the probability of the wavefunctions of the ABSs decay inside the superconductor, as we move away from the junction location.
@@
Let us set up some notations to characterize the S-B-S configuration. We denote the complex superconducting order parameter by $\Delta=\Delta_{0}\,e^{i\,\varphi}$, with $\varphi$ representing the phase variable. Denoting the energy of the eigenstates by $\varepsilon$, a set of discrete states are obtained for $|\varepsilon|<\Delta_{0}$, which are the ABSs refereed to in the discussions above. They are also known as the subgap excitations because their energy values lie within the superconducting gap region. The eigenstates with $|\varepsilon|>\Delta_{0}$ form a continuum, which do not decay within the bulk of the superconductors. Because three bands cross at a single nodal point of a triple-point semimetal, the nature of the ABSs must be different from those arising in the Dirac [34, 36], semi-Dirac [40], Weyl [38], multi-Weyl [38, 39], and RSW [30] semimetals. In particular, for doubly-degenerate nodal points, when the propagation direction is along an axis where dispersion is linear-in-momentum, it has been found that [34, 36, 38] that the energy of the ABSs in the thin-barrier-limit is given by $\varepsilon=\pm\,\Delta_{0}\,\sqrt{1-T_{N}\,\sin^{2}\left(\varphi_{12}/2\right)}$. Here, $\varphi_{12}$ is the difference of the superconducting phases on the two sides of the barrier region [cf. Fig. 1(b)], and $T_{N}$ is the transmission coefficient in an analogous set-up with the two superconducting regions replaced by the normal state of the semimetal. This simple relation arises from the fact that the solution for $\beta\equiv\varepsilon/\Delta_{0}$ is obtained from a polynomial equation of the form ~$\tilde{\mathcal{B}}\cos(2\beta)+\tilde{\mathcal{C}}\cos\varphi_{12}=0$ [36], where $\tilde{\mathcal{B}}$ and $\tilde{\mathcal{C}}$ are purely functions of the barrier strength $V_{0}$, the Fermi energy $E_{F}$, and the magnitude of the perpendicular-to-propagation momentum component $k_{\perp}$. However, such a simple equation does not arise for generic cases, as exemplified by earlier studies on semi-Dirac [40] and RSW [30] semimetals.
@@
Figure 1: Schematics of the (a) dispersion aound a triply-degenerate nodal point; (b) S-B-S junction configuration. For the psedispin-1 semimetal, two of the bands have isotropic linear-in-momentum dispersion behaviour, while the third one is a flat band with momentum-independent constant energy.
The paper is organized as follows. In Sec. II, we consider the 2d triple-point fermions and derive the discrete spectrum for the ABSs resulting from an S-B-S configuration. Therein, we explain how the S-B-S junction is constructed, for theoretical analysis, via the the electron-like and hole-like wavefunctions and the Bogoliubov–de Gennes (BdG) Hamiltonian. This is followed by Sec. III, where we repeat the same exercise for the 3d version of the triple-point fermions. Finally, we end with a summary and an outlook in Sec. IV.
@@
Here, $\mathbf{k}=k_{x}\,\hat{\mathbf{x}}+k_{y}\,\hat{\mathbf{y}}$, and $\bm{\mathcal{S}}$ represents the vector spin-1 operator with the three components
@@
The current density operator for this system is captured by $\check{\mathbf{j}}_{2d}(\mathbf{k})=\nabla_{\mathbf{k}}\mathcal{H}_{2d}(\mathbf{k})=\mathcal{S}_{x}\,\hat{\mathbf{x}}+\mathcal{S}_{y}\,\hat{\mathbf{y}}$. Therefore, for a wavefunction $\psi\equiv\left(c_{1}\quad c_{2}\quad c_{3}\right)^{T}$, we have
@@
This immediately tells us that the contribution of the flat-band wavefunction to the local current density is $\Psi_{0}^{\dagger}\,\check{\mathbf{j}}\,\Psi_{0}=\mathbf{0}\,,$ showing that it does not contribute to transport [18, 19]. Hence, we need to consider only $\Psi_{\pm}$ for computing transport properties.
Eq. (5) also tells us that if the quasiparticles encounter a perpendicular barrier, say at $x=x_{0}$, while propagating along the $x$-direction, the conservation of the probability current density leads to the two boundary conditions [16, 28]
$\lim_{\delta\rightarrow 0}\left[c_{1}(x_{0}-\delta)+c_{3}(x_{0}-\delta)\right]=\lim_{\delta\rightarrow 0}\left[c_{1}(x_{0}+\delta)+c_{3}(x_{0}+\delta)\right]\text{ and }\lim_{\delta\rightarrow 0}c_{2}(x_{0}-\delta)=\lim_{\delta\rightarrow 0}c_{2}(x_{0}+\delta)\,,$ (6)
which we will employ in our derivations. If the Fermi energy cuts the dispersion profile at an energy value $E$, then, for propagation along the $x$-direction, the wavefunction will have the exponential factor $e^{i\,\text{sgn}(E)\,k_{x}\,x}$, where $k_{x}=\sqrt{E^{2}-k_{\perp}^{2}}$ and $k_{\perp}=|k_{y}|$.
II.1 S-B-S junctions oriented perpendicular to the $x$-direction
Here we consider the propagation of the quasiparticles/quasiholes along the $x$-axis. Setting up the S-B-S configuration as shown in Fig. 1(b)], we model the superconducting pair potential as
@@ II.1 S-B-S junctions oriented perpendicular to the $x$-direction
$H_{\text{BdG}}(\mathbf{k})=\begin{pmatrix}\mathcal{H}_{2d}(\mathbf{k})-E_{F}+V(x)&\Delta(x)\\ \Delta^{\dagger}(x)&E_{F}-V(x)-\mathcal{H}_{2d}^{T}(-\mathbf{k})\\ \end{pmatrix},$ (9)
where the subscripts $\{1,2,3\}$ on the fermionic creation and annihilation operators represent the three distinct band indices. We will work assuming the energy-scale hierarchies $V_{0}\gg E_{F}\gg\Delta_{0}$ and $(V_{0}-E_{F})\gg E_{F}$. While the condition $\Delta_{0}\ll E_{F}$ ensures that the mean-field approximation, applicable for using the BdG formalism, is valid, the second condition $(V_{0}-E_{F})\gg E_{F}$ results from taking the thin-barrier limit.
@@
expressed in a piecewise manner for the three regions, where we set the the Fermi energy at $E_{F}$ for the corresponding normal states (i.e., for $\Delta_{0}=0$) in the regions I and III.
@@
    In the left superconductor region, the linear combination, describing the the wavefunction localizing at the interface, must be expressed as
@@
Since the final results depend on the phase difference $\varphi_{12}=\varphi_{2}-\varphi_{1}$, for simplification of the notations, we can set $\varphi_{1}=0$ and $\varphi_{2}=\varphi_{12}$, without any loss of generality. Imposing the continuity of the probability current density at the junctions located at $x=0$ and $x=L$, we get the following conditions [cf. Eq. (6)]:
@@
$\left[\psi_{II}(L,y,k_{\perp})\right]_{1,4}+\left[\psi_{II}(L,y,k_{\perp})\right]_{1,6}=\left[\psi_{III}(L,y,k_{\perp})\right]_{1,4}+\left[\psi_{III}(L,y,k_{\perp})\right]_{1,6},\quad\left[\psi_{II}(L,y,k_{\perp})\right]_{1,5}=\left[\psi_{III}(L,y,L,k_{\perp})\right]_{1,5},$

### v2

@@ Abstract
We study superconductor-barrier-superconductor (S-B-S) Josephson junctions constructed out of two-dimensional and three-dimensional triple-point semimetals, which feature a threefold degeneracy at a single nodal point. We assume a weak and homogeneous s-wave pairing in each superconducting region, and a potential difference is applied across a piece of normal-state semimetal to create the barrier region. We compute the wavefunctions of the Andreev bound states (ABSs), considering the thin-barrier limit. The appropriate boundary conditions at the S-B and B-S junctions allow us to compute the discrete energy eigenvalues ±$\pm|\varepsilon|$ of the ABSs. We get two distinct solutions for $|\varepsilon|$. This result differs from that in graphene and Weyl semimetals, where one obtains only one solution for $|\varepsilon|$. The multifold nature of the triple-point fermions is responsible for this difference. We also illustrate the behaviour of the Josephson current flowing across the S-B-S junction.
Andreev bound states; Josephson junction; pseudospin-1 semimetals
@@ I Introduction
Since the last decade, there have been intensive studies of bandstructures harbouring symmetry-protected band-crossing points in the Brillouin zone (BZ) [1, 2, 3, 4, 5], giving rise to semimetals. The low-energy $\mathbf{k}\cdot\mathbf{p}$ Hamiltonian near the band-crossings gives rise to quasiparticles carrying pseudospin values equal to $\varsigma$, when $(2\,\varsigma+1)$ bands touch at the degeneracy point. We note that, whereas in high-energy physics, a relativistic electron is described by a spin-1/2 fermion (with the intrinsic spin quantum number fixed by Lorentz invariance), the pseudospin of an emergent electronic excitation (still having the spin-1/2 value) in nonrelativistic condensed matter physics can be different from 1/2. In fact, for each of the 230 space groups, the pseudospin quantum numbers are dictated by the irreducible representations of the little group of lattice symmetries at the high-symmetry points in the BZ [3]. The dimensionality of the irreducible representation corresponds to the number of bands crossing at the high-symmetry point. In particular, a threefold degeneracy can arise in a crystal lattice when the nodal point is protected by (I) nonsymmorphic symmetries [3] or (II) symmorphic rotation, combined with mirror symmetries [6, 7, 8, 9]. A schematic illustration of the triply-degenerate nodal point is shown in Fig. 1(a).
@@
In this paper, we focus on triple-point fermionic quasiparticles carrying pseudospin value $\varsigma=1$, whose effective low-energy continuum Hamiltonian is of the form $\mathbf{k}\cdot\bm{\mathcal{S}}$, where $\bm{\mathcal{S}}$ represents the vector of the three matrices which form the spin-1 representation of the SO(3) group. This leads to the emergence of two-dimensional (2d) and three-dimensional (3d) semimetals with pseudospin-1 quasiparticles [10, 11, 12, 5, 13, 14, 15, 6, 7, 16, 8, 17, 18, 19, 20, 21, 22, 23, 24], sometimes dubbed as “Maxwell fermions” [12], since their pseudospin quantum number is analogous to the spin-1 quantum number of the photons (which are described by the Maxwell equations). The 3d versions with linear and isotropic dispersion are the straightforward generalizations of the Weyl semimetal Hamiltonian $\mathbf{k}\cdot\bm{\sigma}$, with $\bm{\sigma}$ representing the vector operator consisting of the three Pauli matrices, thus representing the spin-1/2 representation of the SO(3) group. One can show that the threefold degeneracies carry nonnzero Chern numbers (equalling ±$\pm 2$) [3, 25, 24, 26], analogous to the pseudospin-3/2 Rarita-Schwinger-Weyl (RSW) semimetals [27, 25, 28, 29, 30, 31, 22, 24, 26, 32, 33] harbouring fourfold-degenerate nodal points, thus reflecting the nontrivial topological character of the corresponding bandstructures. Here, we continue the ongoing efforts to unravel the distinct signatures of the triple-point quasiparticles via various transport properties [17, 18, 22, 23, 24]. In particular, the transport property under consideration will be the Josephson current $I_{J}$ [34, 35, 36], arising in configurations consisting of junctions between the normal (abbreviated by “N”) and the superconducting (abbreviated by “S”) phases of 2d and 3d pseudospin-1 semimetals.
@@
When two superconducting regions are coupled to each other by a weak link between them, $I_{J}$ is the equilibrium dissipationless current flowing across the junction. Since the Andreev surface states of the two superconductors hybridize to form Andreev bound states (ABSs) at the junction, they contribute to $I_{J}$. Here, we extend the study of such Josephson effects to 2d and 3d semimetals [37, 38, 39, 40, 41, 42, 33], where the superconducting regions consist of homogeneous s-wave pairing, and the weak link is created by introducing a tunneling barrier. One can consider two alternative arrangements: (I) S-N-S junctions [37, 38]; (II) S-B-S (where “B” indicates a rectangular potential barrier in the N region) junctions [39, 40, 41, 42, 33, 43]. In experiments, superconductivity can be induced in the appropriate region via proximity-effect by placing a conventional s-wave superconductor on top of an electrode made of the semimetal [44], and the barrier region can be created by applying a gate voltage $V_{0}$ across N. The schematics of an S-B-S set-up is illustrated in Fig. 1(b), where we demarcate the left superconducting region as “region I”, the middle barrier region as “region II”, and the right superconducting region as “region III”. We will focus on the short-barrier regime, which implies that the barrier thickness $L$ along the propagation direction is taken to be $L\ll\xi$, where $\xi$ is the superconducting coherence length. In other words, $\xi$ denotes the length scale over which the probability of the wavefunctions of the ABSs decay inside the superconductor, as we move away from the junction location.
@@
Let us set up some notations to characterize the S-B-S configuration. We denote the complex superconducting order parameter by $\Delta=\Delta_{0}\,e^{i\,\varphi}$, with $\varphi$ representing the phase variable. Denoting the energy of the eigenstates by $\varepsilon$, a set of discrete states is obtained for $|\varepsilon|<\Delta_{0}$, which comprises the ABSs referred to in the discussions above. They are also known as the subgap excitations because their energy values lie within the superconducting gap region. The eigenstates with $|\varepsilon|>\Delta_{0}$ form a continuum, which do not decay within the bulk of the superconductors. Because three bands cross at a single nodal point of a triple-point semimetal, the nature of the ABSs must be different from those arising in the Dirac [37, 39], semi-Dirac [43], Weyl [41], multi-Weyl [41, 42], and RSW [33] semimetals. In particular, for doubly-degenerate nodal points, when the propagation direction is along an axis where dispersion is linear-in-momentum, it has been found that [37, 39, 41] that the energy of the ABSs in the thin-barrier-limit is given by $\varepsilon=\pm\,\Delta_{0}\,\sqrt{1-T_{N}\,\sin^{2}\left(\varphi_{12}/2\right)}$. Here, $\varphi_{12}$ is the difference of the superconducting phases on the two sides of the barrier region [cf. Fig. 1(b)], and $T_{N}$ is the transmission coefficient in an analogous set-up with the two superconducting regions replaced by the normal state of the semimetal. This simple relation arises from the fact that the solution for $\beta\equiv\varepsilon/\Delta_{0}$ is obtained from a polynomial equation of the form ~$\tilde{\mathcal{B}}\cos(2\beta)+\tilde{\mathcal{C}}\cos\varphi_{12}=0$ [39], where $\tilde{\mathcal{B}}$ and $\tilde{\mathcal{C}}$ are purely functions of the barrier strength $V_{0}$, the Fermi energy $E_{F}$, and the magnitude of the perpendicular-to-propagation momentum-component, $k_{\perp}$. However, such a simple equation does not arise for generic cases, as exemplified by earlier studies on semi-Dirac [43] and RSW [33] semimetals.
@@
Figure 1: Schematics of the (a) dispersion aound a triply-degenerate nodal point; (b) S-B-S junction configuration. For the pseudospin-1 semimetal, two of the bands demonstrate an isotropic linear-in-momentum dispersion behaviour, while the third one is a flat band with momentum-independent constant energy.
The paper is organized as follows. In Sec. II, we consider the 2d triple-point fermions and derive the discrete spectrum for the ABSs resulting from an S-B-S configuration. Therein, we explain how the S-B-S junction is constructed, for theoretical analysis, via the electron-like and hole-like wavefunctions and the Bogoliubov–de Gennes (BdG) Hamiltonian. This is followed by Sec. III, where we repeat the same exercise for the 3d version of the triple-point fermions. Finally, we end with a summary and an outlook in Sec. IV.
@@
Here, $\mathbf{k}=k_{x}\,\hat{\mathbf{x}}+k_{y}\,\hat{\mathbf{y}}$, $\bm{\mathcal{S}}$ represents the vector spin-1 operator with the three components
@@
The probability-current-density operator for this system is captured by $\check{\mathbf{j}}_{2d}(\mathbf{k})=\nabla_{\mathbf{k}}\mathcal{H}_{2d}(\mathbf{k})=\mathcal{S}_{x}\,\hat{\mathbf{x}}+\mathcal{S}_{y}\,\hat{\mathbf{y}}$. Therefore, for a wavefunction $\psi\equiv\left(c_{1}\quad c_{2}\quad c_{3}\right)^{T}$, the expectation value of the probability flux is given by
@@
This immediately tells us that the contribution of the flat-band wavefunction to the local probability-current density is $\Psi_{0}^{\dagger}\,\check{\mathbf{j}}_{2d}\,\Psi_{0}=\mathbf{0}\,,$ showing that it does not contribute to the determination of the boundary conditions [21, 22]. [^1] See the appendix of Ref. [32] for more details. Although we have the question of propagation of particles only in Ref. [32], and not holes, the discussion equally applies here where we need to consider both the particle- and hole-wavefunctions. Hence, we need to consider only $\Psi_{\pm}$ for imposing the appropriate boundary conditions for a system comprising the normal state. Eq. (5) also tells us that if the quasiparticles encounter a perpendicular barrier, say at $x=x_{0}$, while propagating along the $x$-direction, the conservation of the probability-current density leads to the two boundary conditions [19, 31],
$\lim_{\delta\rightarrow 0}\left[c_{1}(x_{0}-\delta,y)+c_{3}(x_{0}-\delta,y)\right]=\lim_{\delta\rightarrow 0}\left[c_{1}(x_{0}+\delta,y)+c_{3}(x_{0}+\delta,y)\right]\text{ and }\lim_{\delta\rightarrow 0}c_{2}(x_{0}-\delta,y)=\lim_{\delta\rightarrow 0}c_{2}(x_{0}+\delta,y)\,.$ (6)
If the Fermi energy cuts the dispersion profile at an energy value $E$, then for propagation along the $x$-direction, the wavefunction will have the exponential factor $e^{i\,\text{sgn}(E)\,k_{x}\,x}$, where $k_{x}=\sqrt{E^{2}-k_{\perp}^{2}}$ and $k_{\perp}=|k_{y}|$.
II.1 S-B-S junction’s interfaces oriented perpendicular to the $x$-direction
Here we consider the propagation of the quasiparticles/quasiholes along the $x$-axis. We note that the final results would be the same if we had taken the propagation direction to be along the $y$-axis, which follows from the isotropic nature of the dispersion.
Setting up the S-B-S configuration as shown in Fig. 1(b)], we model the superconducting pair potential as
@@ II.1 S-B-S junction’s interfaces oriented perpendicular to the $x$-direction
$H(\mathbf{r})=\frac{1}{2}\sum_{\mathbf{k}}\Psi^{\dagger}_{\mathbf{k}}\,H_{\text{BdG}}(\mathbf{k})\Psi_{\mathbf{k}},\quad\Psi_{\mathbf{k}}=\begin{pmatrix}c_{1}(\mathbf{k})&c_{2}(\mathbf{k})&c_{3}(\mathbf{k})&c_{1}^{\dagger}(-\mathbf{k})&c_{2}^{\dagger}(-\mathbf{k})&c_{3}^{\dagger}(-\mathbf{k})\end{pmatrix}^{T}\,,$
$H_{\text{BdG}}(\mathbf{k},\mathbf{r})=\begin{pmatrix}\mathcal{H}_{2d}(\mathbf{k})-E_{F}+V(x)&\Delta(x)\\ \Delta^{\dagger}(x)&E_{F}-V(x)-\mathcal{H}_{2d}^{T}(-\mathbf{k},\mathbf{r})\\ \end{pmatrix},$ (9)
where the subscripts $\{1,2,3\}$ on the fermionic creation and annihilation operators represent the three distinct band indices. We will work assuming the energy-scale hierarchies embodied by $V_{0}\gg E_{F}\gg\Delta_{0}$ and $(V_{0}-E_{F})\gg E_{F}$. While the condition $\Delta_{0}\ll E_{F}$ ensures that the mean-field approximation, applicable for using the BdG formalism, is valid, the second condition $(V_{0}-E_{F})\gg E_{F}$ results from taking the thin-barrier limit.
@@
$H_{\text{BdG}}(\mathbf{k}\rightarrow-i\bm{\nabla}_{\mathbf{r}},\mathbf{r})\,\psi_{\mathbf{k}}(\mathbf{r})=\varepsilon\,\psi_{\mathbf{k}}(\mathbf{r})\,,$ (10)
@@
Andreev reflection at a normal-metal–superconductor interface couples the electron-like and hole-like wavefunctions. Here we follow the approach outlined in Ref. [37], where the quasiparticles and and the quasiholes are coupled locally by means of a boundary condition on the wavefunction in the normal region. Now, the $x$-component of the probability-current density operator must be modified to
$\check{\mathbf{j}}^{N}_{x}=\partial_{k_{x}}\mathcal{H}_{\text{BdG}}(\mathbf{k},\mathbf{r})=\mathcal{S}_{x}\otimes{\mathcal{S}}_{0}-{\mathcal{S}}_{0}\otimes\mathcal{S}_{x}\,.$ (14)
Therefore, for a wavefunction of the form $\psi_{\text{BdG},0}\equiv\left(c_{1}\quad 0\quad c_{3}\quad c_{4}\quad 0\quad c_{6}\right)^{T}$, we have $\psi_{\text{BdG},0}^{\dagger}\,\check{\mathbf{j}}^{N}_{x}\,\psi_{\text{BdG},0}=0$. Analogous to the discussions of boundary conditions for purely normal-state junctions, we here infer that the electron- and hole-wavefunctions, corresponding to the flat-band state, drop out from the equations arising out of the boundary conditions. Furthermore, Eq. (14) translates to the following constraints, useful for relating the components from a piecewise-defined wavefunction, $\psi_{\text{BdG}}\equiv\left(c_{1}\quad c_{2}\quad c_{3}\quad c_{4}\quad c_{5}\quad c_{6}\right)^{T}$, across an interface at $x=x_{0}$:
$\lim_{\delta\rightarrow 0}\left[c_{1}(x_{0}-\delta,y)+c_{3}(x_{0}-\delta,y)\right]=\lim_{\delta\rightarrow 0}\left[c_{1}(x_{0}+\delta,y)+c_{3}(x_{0}+\delta,y)\right],\lim_{\delta\rightarrow 0}c_{2}(x_{0}-\delta,y)=\lim_{\delta\rightarrow 0}c_{2}(x_{0}+\delta,y)\,.$
$\lim_{\delta\rightarrow 0}\left[c_{4}(x_{0}-\delta)+c_{6}(x_{0}-\delta,y)\right]=\lim_{\delta\rightarrow 0}\left[c_{4}(x_{0}+\delta,y)+c_{6}(x_{0}+\delta,y)\right],\lim_{\delta\rightarrow 0}c_{5}(x_{0}-\delta,y)=\lim_{\delta\rightarrow 0}c_{5}(x_{0}+\delta,y)\,.$ (15)
Here, we have two boundaries: at $x_{0}=0$ and $x_{0}=L$.
@@
expressed in a piecewise manner for the three regions, where we set the Fermi energy at $E_{F}$ for the corresponding normal states (i.e., for $\Delta_{0}=0$) in the regions I and III.
@@
    In the left superconductor region, the linear combination, describing the wavefunction localizing at the interface, must be expressed as
@@
Since the final results depend on the phase difference $\varphi_{12}\equiv\varphi_{2}-\varphi_{1}$, for simplification of the notations, we can set $\varphi_{1}=0$ and $\varphi_{2}=\varphi_{12}$, without any loss of generality. Imposing the continuity of the probability current density at the junctions located at $x=0$ and $x=L$, we get the following conditions [cf. Eq. (II.1)]:
@@
$\left[\psi_{II}(L,y,k_{\perp})\right]_{1,4}+\left[\psi_{II}(L,y,k_{\perp})\right]_{1,6}=\left[\psi_{III}(L,y,k_{\perp})\right]_{1,4}+\left[\psi_{III}(L,y,k_{\perp})\right]_{1,6},\quad\left[\psi_{II}(L,y,k_{\perp})\right]_{1,5}=\left[\psi_{III}(L,y,L,k_{\perp})\right]_{1,5}.$